Olive Company makes silver belt buckles. The company’s master budget appears in the first column of the table.


Required:
Complete the table by preparing Olive’s flexible budget for 4,000, 6,000, and 7,000 units.

Master Budget Flexible Budget Flexible Budget Flexible Budget (5,000 Units) (4,000 Units) (6,000 Units) (7,000 Units) Direct materials Direct labor Variable manufacturing overhead Fixed manufacturing overhead 15,000 S 30,000 S 8,000 S 18,000 S 71,000 $ 12,000 24,000 6,400 18,000 60,400ト Total manufacturing cost
Answer #1

Direct material cost/unit=(15000/5000)=$3

Direct labor cost/unit=(30000/5000)=$6

Variable manufacturing overhead cost/unit=(8000/5000)=$1.6

5000 4000 6000 7000
Direct materials 15000 12000 18000 21000
Direct labor 30000 24000 36000 42000
Variable manufacturing overhead 8000 6400 9600 11200
Fixed manufacturing overhead 18000 18000 18000 18000
Total manufacturing oost 71000 60400 81600 92200

Note:Total fixed costs and variable cost/unit does not change with change in units.